OpenStudy (jojo12g):

Solve the word problem. Luke mowed a lawn today to earn some extra money. He spent $2.10 on fuel and $4.45 on lunch. He charged $13.50 to mow the lawn. What was his profit for mowing the lawn? A. $20.05 B. $15.85 C. $11.15 D. $6.95 @Ms-Brains

OpenStudy (anonymous):

D

OpenStudy (anonymous):

anything else?

OpenStudy (mrcoolguy):

D

MsBrains (ms-brains):

D

MsBrains (ms-brains):

THANK YOU FOR THE MEDAL!

OpenStudy (anonymous):

Solve the word problem. Luke mowed a lawn today to earn some extra money. He spent $2.10 on fuel and $4.45 on lunch. He charged $13.50 to mow the lawn. What was his profit for mowing the lawn? Profit = Gain - Loss P = 13.5 - (2.1 + 4.45) P = 13.5 - 6.55 P = 6.95 D

OpenStudy (jojo12g):

thank u all i appreciate it so muchh thank u thank u thank u

OpenStudy (anonymous):

medal ?

MsBrains (ms-brains):

Welcome. Thanks @madison_taylor914